# A Survey on Occupancy Perception for Autonomous Driving: The Information Fusion Perspective

### v2

@@ 1.3 Contributions
Among perception-related topics, 3D semantic segmentation [16, 17] and 3D object detection [18, 19, 20, 21] have been extensively reviewed. However, these tasks do not facilitate a dense understanding of the environment. BEV perception, which addresses this issue, has also been thoroughly reviewed [2, 1]. Our survey focuses on 3D occupancy perception, which captures the environmental height information overlooked by BEV perception. Roldao et al. [22] conducted a literature review on 3D scene completion for indoor and outdoor scenes, which is close to our focus. Unlike their work, our survey is specifically tailored to autonomous driving scenarios. Moreover, given the multi-source nature of 3D occupancy perception, we provide an in-depth analysis of information fusion techniques for this field. The primary contributions of this survey are three-fold:
@@ 3 Methodologies
Cam4DOcc [10] CVPR 2024 C Vol - 3D Conv F Strong CE - - Cam4DOcc [10]
PanoOcc [4] CVPR 2024 C Vol - MLP, DETR PS Strong Foc, LS - nuScenes [86]
@@
OccGen [95] arXiv 2024 C$+$L Vol - - 3D Conv P Strong CE, LS, Aff - OpenOccupancy [11] - -
Modality: C - Camera; L - LiDAR; R - Radar.
Feature Format: Vol - Volumetric Feature; BEV - Bird’s-Eye View Feature; PV - Perspective View Feature; TPV - Tri-Perspective View Feature.
Lightweight Design: TPV Rp - Tri-Perspective View Representation; Sparse Rp - Sparse Representation.
@@
Task: P - Prediction; F - Forecasting; OP - Open-Vocabulary Prediction; PS - Panoptic Segmentation.
@@
Recent methods of occupancy perception for autonomous driving and their characteristics are detailed in Tab. 1. This table elaborates on the publication venue, input modality, network design, target task, network training and evaluation, and open-source status of each method. Below, we categorize occupancy perception methods into three types, according to the modality of input data. They are LiDAR-centric occupancy perception, vision-centric occupancy perception, and multi-modal occupancy perception. Subsequently, the training of occupancy networks and their loss functions are discussed. Finally, diverse downstream applications that leverage occupancy perception are presented.
@@ 3.2.1 General Pipeline
Inspired by Tesla’s technology of the perception system for their autonomous vehicles [26], vision-centric occupancy perception has garnered significant attention both in industry and academia. Compared to LiDAR-centric methods, vision-centric occupancy perception, which solely relies on camera sensors, represents a current trend. There are three main reasons: (i) Cameras are cost-effective for large-scale deployment on vehicles. (ii) RGB images capture rich environmental textures, aiding in the understanding of scenes and objects such as traffic signs and lane lines. (iii) The burgeoning advancement of deep learning technologies enables a possibility to achieve 3D occupancy perception from 2D vision. Vision-centric occupancy perception can be divided into monocular solutions [102, 53, 25, 50, 51, 32, 54, 88, 84] and multi-camera solutions [52, 103, 30, 37, 60, 79, 104, 31, 81, 8]. Multi-camera perception, which covers a broader field of view, follows a general pipeline as shown in Fig. 5. It begins by extracting front-view feature maps from multi-camera images, followed by a 2D-to-3D transformation, spatial information fusion, and optional temporal information fusion, culminating with an occupancy head that infers the environmental 3D occupancy.
@@ 3.2.3 Information Fusion in Vision-Centric Occupancy
Once the alignment is completed, the historical and current features are fed to the feature fusion module to enhance the representation, especially to strengthen the reasoning ability for occlusion and the recognition ability of moving objects. There are three main streamlines to feature fusion, namely convolution, cross attention, and adaptive mixing. PanoOcc [4] concatenates the previous features with the current ones, then fuses them using a set of 3D residual convolution blocks. Many occupancy perception methods [32, 55, 22, 84, 109] utilize cross attention for fusion. The process is similar to multi-camera cross attention (refer to Eq. 8), but the difference is that 3D-space voxels are projected to 2D multi-frame feature maps instead of multi-camera feature maps. Moreover, FullySparse [107] employs adaptive mixing [115] for temporal information fusion. For the query feature of the current frame, FullySparse samples $S_{n}$ features from historical frames, and aggregates them through adaptive mixing. Specifically, the sampled features are multiplied by the channel mixing matrix $W_{C}$ and the point mixing matrix $W_{S_{n}}$, respectively. These mixing matrices are dynamically generated from the query feature $F_{q}$ of the current frame:
@@ 3.4.1 Training with Strong Supervision
where $\left(V,\hat{V}\right)$ are the ground-truth and predicted semantic occupancy with $N_{C}$ categories. $\omega_{c}$ is a weight for a specific class $c$ according to the inverse of the class frequency. Moreover, some occupancy perception methods use other semantic losses, which are commonly utilized in semantic segmentation tasks [123, 124], such as Lovasz-Softmax loss [125] and Focal loss [126]. Furthermore, there are two specialized semantic losses: frustum proportion loss [25], which provides cues to alleviate occlusion ambiguities from the perspective of the visual frustum, and position awareness loss [127], which leverages local semantic entropy to encourage sharper semantic and geometric gradients.
@@ 4.1.2 Metrics
(2) Ray-level Metric: Although voxel-level IoU and mIoU metrics are widely recognized [10, 85, 79, 52, 37, 87, 75, 84], they still have limitations. Due to unbalanced distribution and occlusion of LiDAR sensing, ground-truth voxel labels from accumulated LiDAR point clouds are imperfect, where the areas not scanned by LiDAR are annotated as empty. Moreover, for thin objects, voxel-level metrics are too strict, as a one-voxel deviation would reduce the IoU values of thin objects to zero. To solve these issues, FullySparse [107] imitates LiDAR’s ray casting and proposes ray-level mIoU, which evaluates rays to their closest contact surface. This novel mIoU, combined with the mean absolute velocity error (mAVE), is adopted by the occupancy score (OccScore) metric [93]. OccScore overcomes the shortcomings of voxel-level metrics while also evaluating the performance in perceiving object motion in the scene (i.e., occupancy flow).
@@ 4.2.1 Perception Accuracy
FB-OCC [140] Strong 42.06 - 14.30 49.71 30.0 46.62 51.54 29.3 29.13 29.35 30.48 34.97 39.36 83.07 47.16 55.62 59.88 44.89 39.58
PanoOcc [4] Strong 42.13 - 11.67 50.48 29.64 49.44 55.52 23.29 33.26 30.55 30.99 34.43 42.57 83.31 44.23 54.4 56.04 45.94 40.4
@@
SemanticKITTI [59] is the first dataset with 3D occupancy labels for outdoor driving scenes. Occ3D-nuScenes [35] is the dataset used in the CVPR 2023 3D Occupancy Prediction Challenge [141]. These two datasets are currently the most popular. Therefore, we summarize the performance of various 3D occupancy methods that are trained and tested on these datasets, as reported in Tab. 3 and 4. These tables further organize occupancy methods according to input modalities and supervised learning types, respectively. The best performances are highlighted in bold. Tab. 3 utilizes the IoU and mIoU metrics to evaluate the 3D geometry and 3D semantics occupancy perception capabilities. Tab. 4 adopts mIoU and mIoU to assess semantic occupancy perception. Unlike mIoU, the mIoU metric excludes ’others’ and ’other flat’ classes and is used by the self-supervised OccNeRF [37]. For fairness, we compute mIoU of other self-supervised occupancy methods. Notably, the OccScore metric is used in the CVPR 2024 Autonomous Grand Challenge [142], but is not universal currently. Thus, we do not summarize the occupancy performance with this metric. Below, we will compare perception accuracy from three aspects: overall comparison, modality comparison, and supervision comparison.
@@
(1) Overall Comparison. Tab. 3 shows that (i) the IoU scores of occupancy networks are less than 50%, while the mIoU scores are less than 30%. The IoU scores (indicating geometric perception, i.e., ignoring semantics) substantially surpass the mIoU scores. This is because predicting occupancy for some semantic categories is challenging, such as bicycles, motorcycles, persons, bicyclists, motorcyclists, poles, and traffic signs. Each of these classes has a small proportion (under 0.3%) in the dataset, and their small sizes in shapes make them difficult to observe and detect. Therefore, if the IOU scores of these categories are low, they significantly impact the overall mIoU value. Because the mIOU calculation, which does not account for category frequency, divides the total IoU scores of all categories by the number of categories. (ii) A higher IoU does not guarantee a higher mIoU. One possible explanation is that the semantic perception capacity (reflected in mIoU) and the geometric perception capacity (reflected in IoU) of an occupancy network are distinct and not positively correlated.
Form Tab. 4, it is evident that (i) the mIOU scores of occupancy networks are within 50%, higher than the scores on SemanticKITTI. For example, the mIOU of TPVFormer [31] on SemanticKITTI is 11.26%, but it gets 27.83% on Occ3D-nuScenes. Similarly, OccFormer [54] and SurroundOcc [75] have the same situation. We consider this might be due to the more accurate occupancy labels in Occ3D-nuScenes. SemanticKITTI annotates each voxel based on the LiDAR point cloud, that is, assigns labels to voxels based on a majority vote over all labelled points inside the voxel. In contrast, Occ3D-nuScenes utilizes a complex label generation process, including voxel densification, occlusion reasoning, and image-guided voxel refinement. This annotation could produce more precise and dense 3D occupancy labels. (ii) COTR [85] has the best mIoU (46.21%) and also achieves the highest scores in IoU across all categories.
@@
(2) Modality Comparison. The input data modality significantly influences on 3D occupancy perception accuracy. The ’Mod.’ column in Tab. 3 reports the input modality of various occupancy methods. It can be seen that, due to the accurate depth information provided by LiDAR sensing, LiDAR-centric occupancy methods have more precise perception with higher IoU and mIoU scores. For example, S3CNet [29] has the top mIoU (29.53%) and DIFs [74] achieves the highest IoU (58.90%). We observe that the two multi-modal approaches do not outperform S3CNet and DIFs, indicating that they have not fully leveraged the benefits of multi-modal fusion and the richness of input data. There is considerable potential for further improvement in multi-modal occupancy perception. Besides, although vision-centric occupancy perception has advanced rapidly in recent years, as can be seen from Tab. 3, state-of-the-art vision-centric occupancy methods still have a gap with LiDAR-centric methods in terms of IoU and mIoU. We consider that further improving depth estimation for vision-centric methods is necessary.

### v3

@@ 1.3 Contributions
Among perception-related topics, 3D semantic segmentation [16, 17] and 3D object detection [18, 19, 20, 21] have been extensively reviewed. However, these tasks do not facilitate a dense understanding of the environment. BEV perception, which addresses this issue, has also been thoroughly reviewed [1, 2]. Our survey focuses on 3D occupancy perception, which captures the environmental height information overlooked by BEV perception. There are two related reviews: Roldao et al. [22] conducted a literature review on 3D scene completion for both indoor and outdoor scenes; Zhang et al. [23] only reviewed 3D occupancy prediction based on the visual modality. Unlike their work, our survey is tailored to autonomous driving scenarios, and extends the existing 3D occupancy survey by considering more sensor modalities. Moreover, given the multi-source nature of 3D occupancy perception, we provide an in-depth analysis of information fusion techniques for this field. The primary contributions of this survey are three-fold:
@@ 3 Methodologies
Cam4DOcc [10] CVPR 2024 C Vol - 3D Conv P&F Strong CE - - Cam4DOcc [10]
PanoOcc [4] CVPR 2024 C Vol - MLP, DETR PO Strong Foc, LS - nuScenes [86]
@@
Bi-SSC [94] CVPR 2024 C BEV - - 3D Conv P Strong CE, Aff - - - -
LowRankOcc [95] CVPR 2024 C Vol - - TRDR Mask Decoder P Strong MC - SurroundOcc [76] - -
PaSCo [96] CVPR 2024 L Vol - - - Mask Decoder PO Strong CE, LS, MC - SSCBench [37], Robo3D [97]
HTCL [98] ECCV 2024 C Vol - 3D Conv P Strong CE, Aff - OpenOccupancy [11]
OSP [99] ECCV 2024 C Point - - MLP P Strong CE - -
OccGen [100] ECCV 2024 C$+$L Vol - - 3D Conv P Strong CE, LS, Aff - OpenOccupancy [11] - -
Scribble2Scene [101] IJCAI 2024 C Vol - - MLP P Weak CE, Aff, KD - SemanticPOSS [102] - -
@@
Modality: C - Camera; L - LiDAR; R - Radar; T - Text.
Feature Format: Vol - Volumetric Feature; BEV - Bird’s-Eye View Feature; PV - Perspective View Feature; TPV - Tri-Perspective View Feature; Point - Point Feature.
Lightweight Design: TPV Rp - Tri-Perspective View Representation; Sparse Rp - Sparse Representation; TRDR - Tensor Residual Decomposition and Recovery.
@@
Task: P - Prediction; F - Forecasting; OP - Open-Vocabulary Prediction; PO - Panoptic Occupancy.
@@
Recent methods of occupancy perception for autonomous driving and their characteristics are detailed in Tab. 1. This table elaborates on the publication venue, input modality, network design, target task, network training and evaluation, and open-source status of each method. In this section, according to the modality of input data, we categorize occupancy perception methods into three types: LiDAR-centric occupancy perception, vision-centric occupancy perception, and multi-modal occupancy perception. Additionally, network training strategies and corresponding loss functions will be discussed.
@@ 3.2.1 General Pipeline
Inspired by Tesla’s technology of the perception system for their autonomous vehicles [27], vision-centric occupancy perception has garnered significant attention both in industry and academia. Compared to LiDAR-centric methods, vision-centric occupancy perception, which solely relies on camera sensors, represents a current trend. There are three main reasons: (i) Cameras are cost-effective for large-scale deployment on vehicles. (ii) RGB images capture rich environmental textures, aiding in the understanding of scenes and objects such as traffic signs and lane lines. (iii) The burgeoning advancement of deep learning technologies enables a possibility to achieve 3D occupancy perception from 2D vision. Vision-centric occupancy perception can be divided into monocular solutions [26, 33, 51, 52, 54, 55, 84, 88, 111, 112] and multi-camera solutions [8, 31, 32, 38, 53, 61, 80, 81, 113, 114, 115]. Multi-camera perception, which covers a broader field of view, follows a general pipeline as shown in Fig. 5. It begins by extracting front-view feature maps from multi-camera images, followed by a 2D-to-3D transformation, spatial information fusion, and optional temporal information fusion, culminating with an occupancy head that infers the environmental 3D occupancy.
@@ 3.2.3 Information Fusion in Vision-Centric Occupancy
Once the alignment is completed, the historical and current features are fed to the feature fusion module to enhance the representation, especially to strengthen the reasoning ability for occlusion and the recognition ability of moving objects. There are three main streamlines to feature fusion, namely convolution, cross attention, and adaptive mixing. PanoOcc [4] concatenates the previous features with the current ones, then fuses them using a set of 3D residual convolution blocks. Many occupancy perception methods [22, 33, 56, 84, 120] utilize cross attention for fusion. The process is similar to multi-camera cross attention (refer to Eq. 8), but the difference is that 3D-space voxels are projected to 2D multi-frame feature maps instead of multi-camera feature maps. Moreover, SparseOcc [118] [^1] Concurrently, two works with the same name SparseOcc [90, 118] explore sparsity in occupancy from different directions. employs adaptive mixing [126] for temporal information fusion. For the query feature of the current frame, SparseOcc samples $S_{n}$ features from historical frames, and aggregates them through adaptive mixing. Specifically, the sampled features are multiplied by the channel mixing matrix $W_{C}$ and the point mixing matrix $W_{S_{n}}$, respectively. These mixing matrices are dynamically generated from the query feature $F_{q}$ of the current frame:
@@ 3.4.1 Training with Strong Supervision
where $\left(V,\hat{V}\right)$ are the ground-truth and predicted semantic occupancy with $N_{C}$ categories. $\omega_{c}$ is a weight for a specific class $c$ according to the inverse of the class frequency. Notably, CE loss and BCE loss are also widely used in semantic segmentation [134, 135]. Besides these losses, some occupancy perception methods employ other semantic losses commonly utilized in semantic segmentation tasks [136, 137], such as Lovasz-Softmax loss [138] and Focal loss [139]. Furthermore, there are two specialized semantic losses: frustum proportion loss [26], which provides cues to alleviate occlusion ambiguities from the perspective of the visual frustum, and position awareness loss [140], which leverages local semantic entropy to encourage sharper semantic and geometric gradients.
@@ 3.4.2 Training with Other Supervisions
Training with strong supervision is straightforward and effective, but requires tedious annotation for voxel-wise labels. In contrast, training with other types of supervision, such as weak, semi, and self supervision, is label-efficient.
@@ 4 Evaluation
In this section, we will provide the performance evaluation of 3D occupancy perception. First, the datasets and metrics commonly used for evaluation are introduced. Subsequently, we offer detailed performance comparisons and discussions on state-of-the-art 3D occupancy perception methods using the most popular datasets.
@@ 4.1.2 Metrics
(2) Ray-level Metric: Although voxel-level IoU and mIoU metrics are widely recognized [10, 38, 53, 76, 80, 84, 85, 87], they still have limitations. Due to unbalanced distribution and occlusion of LiDAR sensing, ground-truth voxel labels from accumulated LiDAR point clouds are imperfect, where the areas not scanned by LiDAR are annotated as empty. Moreover, for thin objects, voxel-level metrics are too strict, as a one-voxel deviation would reduce the IoU values of thin objects to zero. To solve these issues, SparseOcc [118] imitates LiDAR’s ray casting and proposes ray-level mIoU, which evaluates rays to their closest contact surface. This novel mIoU, combined with the mean absolute velocity error (mAVE), is adopted by the occupancy score (OccScore) metric [93]. OccScore overcomes the shortcomings of voxel-level metrics while also evaluating the performance in perceiving object motion in the scene (i.e., occupancy flow).
@@ 4.2 Performance
In this subsection, we will compare and analyze the performance accuracy and inference speed of various 3D occupancy perception methods. For performance accuracy, we discuss three aspects: overall comparison, modality comparison, and supervision comparison. The evaluation datasets used include SemanticKITTI, Occ3D-nuScenes, and SSCBench-KITTI-360.
@@ 4.2.1 Perception Accuracy
FB-OCC [153] Strong 42.06 - 14.30 49.71 30.00 46.62 51.54 29.30 29.13 29.35 30.48 34.97 39.36 83.07 47.16 55.62 59.88 44.89 39.58
PanoOcc [4] Strong 42.13 - 11.67 50.48 29.64 49.44 55.52 23.29 33.26 30.55 30.99 34.43 42.57 83.31 44.23 54.40 56.04 45.94 40.40
@@
Table 5: 3D occupancy benchmarking results (%) on the SSCBench-KITTI-360 test set. The best results are in bold. OccFiner (Mono.) indicates that OccFiner refines the predicted occupancy from MonoScene.
Method IoU mIoU
$\blacksquare$ car (2.85%)
$\blacksquare$ bicycle (0.01%)
$\blacksquare$ motorcycle (0.01%)
$\blacksquare$ truck (0.16%)
$\blacksquare$ other-veh. (5.75%)
$\blacksquare$ person (0.02%)
$\blacksquare$ road (14.98%)
$\blacksquare$ parking (2.31%)
$\blacksquare$ sidewalk (6.43%)
$\blacksquare$ other-grnd. (2.05%)
$\blacksquare$ building (15.67%)
$\blacksquare$ fence (0.96%)
$\blacksquare$ vegetation (41.99%)
$\blacksquare$ terrain (7.10%)
$\blacksquare$ pole (0.22%)
$\blacksquare$ traf.-sign (0.06%)
$\blacksquare$ other-struct. (4.33%)
$\blacksquare$ other-obj. (0.28%)
LiDAR-Centric Methods
SSCNet [25] 53.58 16.95 31.95 0.00 0.17 10.29 0.00 0.07 65.70 17.33 41.24 3.22 44.41 6.77 43.72 28.87 0.78 0.75 8.69 0.67
LMSCNet [28] 47.35 13.65 20.91 0.00 0.00 0.26 0.58 0.00 62.95 13.51 33.51 0.20 43.67 0.33 40.01 26.80 0.00 0.00 3.63 0.00
Vision-Centric Methods
GaussianFormer [154] 35.38 12.92 18.93 1.02 4.62 18.07 7.59 3.35 45.47 10.89 25.03 5.32 28.44 5.68 29.54 8.62 2.99 2.32 9.51 5.14
MonoScene [26] 37.87 12.31 19.34 0.43 0.58 8.02 2.03 0.86 48.35 11.38 28.13 3.32 32.89 3.53 26.15 16.75 6.92 5.67 4.20 3.09
OccFiner (Mono.) [155] 38.51 13.29 20.78 1.08 1.03 9.04 3.58 1.46 53.47 12.55 31.27 4.13 33.75 4.62 26.83 18.67 5.04 4.58 4.05 3.32
VoxFormer [33] 38.76 11.91 17.84 1.16 0.89 4.56 2.06 1.63 47.01 9.67 27.21 2.89 31.18 4.97 28.99 14.69 6.51 6.92 3.79 2.43
TPVFormer [32] 40.22 13.64 21.56 1.09 1.37 8.06 2.57 2.38 52.99 11.99 31.07 3.78 34.83 4.80 30.08 17.52 7.46 5.86 5.48 2.70
OccFormer [55] 40.27 13.81 22.58 0.66 0.26 9.89 3.82 2.77 54.30 13.44 31.53 3.55 36.42 4.80 31.00 19.51 7.77 8.51 6.95 4.60
Symphonies [88] 44.12 18.58 30.02 1.85 5.90 25.07 12.06 8.20 54.94 13.83 32.76 6.93 35.11 8.58 38.33 11.52 14.01 9.57 14.44 11.28
CGFormer [156] 48.07 20.05 29.85 3.42 3.96 17.59 6.79 6.63 63.85 17.15 40.72 5.53 42.73 8.22 38.80 24.94 16.24 17.45 10.18 6.77
@@
SemanticKITTI [60] is the first dataset with 3D occupancy labels for outdoor driving scenes. Occ3D-nuScenes [36] is the dataset used in the CVPR 2023 3D Occupancy Prediction Challenge [157]. These two datasets are currently the most popular. Therefore, we summarize the performance of various 3D occupancy methods that are trained and tested on these datasets, as reported in Tab. 3 and 4. Additionally, we evaluate the performance of 3D occupancy methods on the SSCBench-KITTI-360 dataset, as reported in Tab. 5. These tables classify occupancy methods according to input modalities and supervised learning types, respectively. The best performances are highlighted in bold. Tab. 3 and 5 utilize the IoU and mIoU metrics to evaluate the 3D geometric and 3D semantic occupancy perception capabilities. Tab. 4 adopts mIoU and mIoU to assess 3D semantic occupancy perception. Unlike mIoU, the mIoU metric excludes the ’others’ and ’other flat’ classes and is used by the self-supervised OccNeRF [38]. For fairness, we compare the mIoU of OccNeRF with other self-supervised occupancy methods. Notably, the OccScore metric is used in the CVPR 2024 Autonomous Grand Challenge [158], but it has yet to become widely adopted. Thus, we do not summarize the occupancy performance with this metric. Below, we will compare perception accuracy from three aspects: overall comparison, modality comparison, and supervision comparison.
@@
(1) Overall Comparison. Tab. 3 and 5 show that (i) the IoU scores of occupancy networks are less than 60%, while the mIoU scores are less than 30%. The IoU scores (indicating geometric perception, i.e., ignoring semantics) substantially surpass the mIoU scores. This is because predicting occupancy for some semantic categories is challenging, such as bicycles, motorcycles, persons, bicyclists, motorcyclists, poles, and traffic signs. Each of these classes has a small proportion (under 0.3%) in the dataset, and their small sizes in shapes make them difficult to observe and detect. Therefore, if the IOU scores of these categories are low, they significantly impact the overall mIoU value. Because the mIOU calculation, which does not account for category frequency, divides the total IoU scores of all categories by the number of categories. (ii) A higher IoU does not guarantee a higher mIoU. One possible explanation is that the semantic perception capacity (reflected in mIoU) and the geometric perception capacity (reflected in IoU) of an occupancy network are distinct and not positively correlated.
Form Tab. 4, it is evident that (i) the mIOU scores of occupancy networks are within 50%, higher than the scores on SemanticKITTI and SSCBench-KITTI-360. For example, the mIOUs of TPVFormer [32] on SemanticKITTI and SSCBench-KITTI-360 are 11.26% and 13.64%, but it gets 27.83% on Occ3D-nuScenes. OccFormer [55] and SurroundOcc [76] have similar situations. We consider this might be due to the simpler task setting in Occ3D-nuScenes. On the one hand, Occ3D-nuScenes uses surrounding-view images as input, containing richer scene information compared to SemanticKITTI and SSCBench-KITTI-360, which only utilize monocular or binocular images. On the other hand, Occ3D-nuScenes only calculates mIOU for visible 3D voxels, whereas the other two datasets evaluate both visible and occluded areas, posing greater challenges. (ii) COTR [85] has the best mIoU (46.21%) and also achieves the highest scores in IoU across all categories on Occ3D-nuScenes.
@@
(2) Modality Comparison. The input data modality significantly influences 3D occupancy perception accuracy. Tab. 3 and 5 report the performance of occupancy perception in different modalities. It can be seen that, due to the accurate depth information provided by LiDAR sensing, LiDAR-centric occupancy methods have more precise perception with higher IoU and mIoU scores. For example, on the SemanticKITTI dataset, S3CNet [30] has the top mIoU (29.53%) and DIFs [75] achieves the highest IoU (58.90%); on the SSCBench-KITTI-360 dataset, S3CNet achieves the best IoU (53.58%). However, we observe that the multi-modal approaches (e.g., OpenOccupancy [11] and Co-Occ [103]) do not outperform single-modal (i.e., LiDAR-centric or vision-centric) methods, indicating that they have not fully leveraged the benefits of multi-modal fusion and the richness of input data. Therefore, there is considerable potential for further improvement in multi-modal occupancy perception. Moreover, vision-centric occupancy perception has advanced rapidly in recent years. On the SemanticKITTI dataset, the state-of-the-art vision-centric occupancy methods still lag behind LiDAR-centric methods in terms of IoU and mIoU. But notably, the mIoU of the vision-centric CGFormer [156] has surpassed that of LiDAR-centric methods on the SSCBench-KITTI-360 dataset.